Solve the following problem.


1. Calculate the oH of a butter system containing 1.0 M CH3COOH and 1.0 M CH3COONa.

2. What is the pH of the butter system after the addition of 0.10 mole of gaseous HCI to 1 L of the solution.


1
Expert's answer
2021-06-22T04:42:39-0400

1). I think the m in 1.0 m should actually be M (molar) not molal. Same for the other concentration.

To find the pH of a buffer, it's usually easiest to just use the Henderson Hasselbalch equation:

pH = pKa + log [salt]/[acid]

pKa = -log Ka = -log 1.8x10-5 = 4.74

pH = 4.74 + log (0.50 M/1.0 M) = 4.74 + log 0.5 = 4.74 + (-0.3)

pH = 4.74 - 0.3

pH = 4.44

2). 0.01L × 0.1 mole HCl/1L = 1 × 10-3 mol HCl

pH = -log (1×10-3)

pH = 3